# MOTIF: Multi-strategy Optimization via Turn-based Interactive Framework

### v1

@@ A.4 Self-play and Adversarial Learning in LLMs
Recent work has shown that self-play—where language models act as interacting agents—can lead to significant gains in reasoning quality, robustness, and decision-making. Inspired by game-theoretic dynamics, these methods structure LLM interaction through alternating roles such as attacker–defender, proposer–critic, or buyer–seller, enabling emergent behavior that improves model output beyond static prompting. For instance, SPAG (Cheng et al. 2025) frames reasoning as a two-player adversarial game in which an attacker introduces misleading arguments and a defender must debunk them, improving the model’s capacity for deception detection. Similarly, the CDG framework (Wang et al. 2025) trains a pair of prover and critic agents to iteratively expose and correct reasoning flaws, leading to more logically sound outputs.
Other work explores negotiation and cooperation in self-play. Fu et al. (Fu et al. 2023a) design an in-context feedback loop in which LLMs simulate negotiation roles (e.g., buyer vs. seller) and revise their proposals based on AI-generated critiques, resulting in improved utility and coherence in generated dialogues. These methods typically rely on prompt structuring, role conditioning, and iterative critique to create a feedback-rich interaction landscape—one that mirrors social or argumentative reasoning among humans.
@@
Adversarial learning in LLMs also intersects with verbal reinforcement learning and reflective prompting. For example, Reflexion (Shinn et al. 2023) guides the model to reflect on its own past errors and learn through verbalized feedback, strengthening multi-step reasoning capabilities. Although these studies focus primarily on tasks like logic puzzles, negotiation, and QA, their results suggest that interactive multi-agent dynamics offer a powerful foundation for tasks that benefit from critique, revision, or competition. Taken together, these approaches illustrate how adversarial pressure and structured self-interaction can improve LLM outputs by introducing feedback loops, multiple perspectives, and dynamic behavior adjustment—offering a foundation for more sophisticated systems that go beyond static prompting or single-agent search.
@@ B.1 Benchmark Problems
The Capacitated Vehicle Routing Problem (CVRP) extends the TSP by introducing multiple vehicles that serve customers under capacity constraints. Let $G=(V,E)$ be a complete undirected graph where $V=\{v_{0},v_{1},\dots,v_{n}\}$ consists of a depot node $v_{0}$ and $n$ customer nodes, and let $c(i,j)\in\mathbb{R}_{\geq 0}$ denote the travel cost between nodes $v_{i}$ and $v_{j}$. Each customer $v_{i}$ has a demand $d_{i}>0$, and all vehicles are identical with a fixed capacity $Q$. A route is a sequence $r=(v_{0},v_{i_{1}},\dots,v_{i_{m}},v_{0})$ in which a single vehicle departs from the depot, serves a subset of customers, and returns to the depot. The cost of such a route is given by
@@ B.2 Benchmark Algorithms
Guided Local Search (GLS) is a well-established metaheuristic designed to escape local optima by penalizing frequently occurring features in poor-quality solutions. Introduced by Voudouris and Tsang (1999), GLS enhances standard local search by augmenting the objective function with feature-based penalties, thereby encouraging diversification while maintaining efficiency. It has demonstrated strong performance on various combinatorial problems, particularly the TSP, and has been used in hybrid solvers such as GLS-PR for Vehicle Routing Problems (VRP) (Shaw 1997, 1998) and more recently in neural-guided contexts (Hudson et al. 2022; Sui et al. 2024).
@@
2:Main strategy: $G\leftarrow$ GenerateGuideMatrix($D$)
@@
[…]
9:for $t=1$ to $T$ do
10:  for $m=1$ to $M$ do
11:   $(i^{*},j^{*})\leftarrow\arg\max_{(i,j)\in\sigma_{\text{cur}}}\frac{G_{ij}}{1+P_{ij}}$
12:   $P_{i^{*}j^{*}}\leftarrow P_{i^{*}j^{*}}+1$
13:   $\hat{D}\leftarrow D+k\cdot P$
[…]
15:  end for
[…]
17:  if $\textsc{Cost}(D,\sigma_{\text{cur}})<c_{\text{best}}$ then
18:   $\sigma_{\text{best}}\leftarrow\sigma_{\text{cur}}$
19:   $c_{\text{best}}\leftarrow\textsc{Cost}(D,\sigma_{\text{cur}})$
20:  end if
21:end for
22:Return: $\sigma_{\text{best}}$
@@
Ant Colony Optimization (ACO) is a metaheuristic inspired by the foraging behavior of real ants, first introduced by Dorigo, Maniezzo, and Colorni (1996) as the Ant System for solving COP s. In ACO, a colony of artificial ants incrementally construct solutions by moving on a problem graph, guided by probabilistic rules that balance pheromone intensity and heuristic information. After completing a solution, each ant deposits pheromone on visited components, reinforcing promising paths over iterations. The Ant Colony System (ACS) (Dorigo and Gambardella 1997) refined this process by introducing local pheromone updates and elitist reinforcement strategies.
Since its inception, ACO has been extended in various directions, including hybridization with local search, parallel implementations, and learning-based variants. Recent approaches such as DeepACO (Ye et al. 2023) incorporate neural networks to guide construction policies or predict pheromone distributions, significantly improving scalability and adaptability to different problem instances. These developments have established ACO as a versatile and extensible baseline for many routing and packing tasks.
We now formalize the ACO solver structure used across our benchmark tasks. Each solver consists of three strategies—initialization, transition rule, and pheromone update—optimized independently. We provide a unified pseudocode and detail problem-specific implementations in Table 5.
@@
While prior work in AHD typically focuses on optimizing a single heuristic matrix $H$ (e.g., based on inverse distances), our formulation broadens this space in several ways. First, we additionally treat the initialization of the pheromone matrix $P$ as part of the optimization process. Second, we model the combination of $H$ and $P$ into transition weights as a learnable strategy, rather than using a fixed formula. Third, we expose the pheromone update rule itself to adaptation. Finally, we extend the function signatures to include both the current iteration $t$ and the total horizon $T$, enabling LLMs to discover temporally adaptive behaviors that prior formulations cannot express.
@@
The idea of deconstruction and repair is not new in combinatorial optimization. André and Kevin (2020) proposed a neural variant of Large Neighborhood Search (LNS), where destroy and repair policies are jointly learned via reinforcement learning. Wu et al. (2021) further improved this framework by leveraging GNN-based representations for adaptive neighborhood selection. More recently, Fu et al. (2023b) introduced a hierarchical destroy-and-repair framework that scales to million-city TSP instances by progressively refining solutions across multiple levels. In parallel, Li et al. (2025) explored hypergraph-based operators to guide destruction and reconstruction phases, achieving state-of-the-art results.
@@ C.1 Prompt Structure
The system prompt is a persistent instruction that defines the identity and behavior of the lLLM across interactions. Unlike user inputs, which vary per query, it provides a stable context that aligns the model’s responses with task objectives. A common technique is role-based prompting, where assigning the LLM a specific persona—e.g., “competitive algorithm designer”—helps activate relevant reasoning styles. This role-play approach has been shown to enhance zero-shot performance by encouraging coherence, structured thinking, and task relevance (Kong et al. 2024).
@@ 3. Ensure code is syntactically correct and handles edge cases
Implement the initialization strategy that sets up guidance matrices for route finding:.
1import numpy as np 2 3def initialize(distances: np.ndarray) -> tuple[np.ndarray, np.ndarray]: 4 """ 5 Initialize heuristic and pheromone matrices for route optimization. 6 7 distances : np.ndarray, shape (n_cities, n_cities) 8 heuristic : np.ndarray, shape (n_cities, n_cities) 9 Matrix representing desirability of traveling between cities 10 pheromone : np.ndarray, shape (n_cities, n_cities) 11 Matrix representing initial intensity of guidance trails 12 """ 13 # Your implementation here 14 pass <</task>>
@@
The human message, framed from a third-person referee perspective, delivers the competition status to the agent while simultaneously acting as a coach—offering guidance and exerting competitive pressure. Notably, we intentionally place the operator instructions at the end of the prompt to draw the model’s attention, leveraging findings from Liu et al. (2023), which show that LLMs often overlook information positioned in the middle of long prompts.
@@ C.3 First Round: Component-wise Competition
1:Input: Initial strategy set $\boldsymbol{\Pi}=(\pi_{1},\ldots,\pi_{K})$, outer iterations $T_{\text{outer}}$, inner iterations $T_{\text{inner}}$
@@
3:for $k=1$ to $K$ do
4:  $\mathcal{T}_{k}\leftarrow\textsc{InitTree}(\pi_{k},C_{0})$
5:  $N_{k}\leftarrow 0\quad$ {Visit count}
6:  $R_{k}\leftarrow 0\quad$ {Cumulative reward}
7:end for
8:for $t=1$ to $T_{\text{outer}}$ do
9:  Select strategy index:
$k^{\ast}\leftarrow\arg\max_{1\leq k\leq K}\left(\frac{R_{k}}{N_{k}}+C_{\text{out}}\cdot\sqrt{\frac{\ln(\sum_{j}N_{j}+1)}{N_{k}}}\right)$
10:  $\pi_{k^{\ast}}^{\prime}\leftarrow\textsc{CompetitiveMCTS}(\mathcal{T}_{k^{\ast}},T_{\text{inner}})\quad$ {Algorithm 6}
11:  $\boldsymbol{\Pi}[k^{\ast}]\leftarrow\pi_{k^{\ast}}^{\prime}$
12:  $C_{t}\leftarrow\textsc{Evaluate}(\boldsymbol{\Pi})$
13:  if $C_{t}<C_{0}$ then
14:   $I\,(\%)\leftarrow(C_{0}-C_{t})/|C_{0}|\quad$ {II: improvement over baseline}
15:   $R_{k^{\ast}}\leftarrow R_{k^{\ast}}+\sigma(I)\quad$ $\sigma$: sigmoid function}
16:   $C_{0}\leftarrow C_{t}$
17:   Propagate baseline $C_{0}$ to all $\mathcal{T}_{k}\quad$ {dynamic baseline}
18:  else
19:   $R_{k^{\ast}}\leftarrow R_{k^{\ast}}+0.5$
20:   Revert $\pi_{k^{\ast}}$ to previous best
21:  end if
22:  $N_{k^{\ast}}\leftarrow N_{k^{\ast}}+1$
23:end for
24:Return: Best strategies $\boldsymbol{\Pi}$
@@
where $F(s\!\to\!s^{\prime})=U(s^{\prime})-U(s)$ is exactly in the form of a potential‐based shaping reward ($\Phi(s^{\prime})-\Phi(s)$), which by Ng, Harada, and Russell (1999) guarantees policy invariance. $G(s)=U(s)$ depends only on the current state $s$ (i.e. is constant across all child actions), and thus does not affect the relative ranking of successor Q‐values in MCTS selection.
@@
1:Input: Root node $n_{0}$ storing $\pi_{1}(n_{0}),\,\pi_{2}(n_{0})$, baseline cost $C_{0}$, operator set $\mathcal{O}$, number of iterations $T$
2: Initialize $C^{*}_{1},C^{*}_{2}\leftarrow\infty$, $\pi^{*}_{1}\leftarrow\pi_{1}(n_{0})$, $\pi^{*}_{2}\leftarrow\pi_{2}(n_{0})$, current player $p\leftarrow 1$
@@
4:  $n\leftarrow n_{0}$ 1. SELECTION:
5:  while there is at least one $o\in\mathcal{O}$ without a child annotated $(o,p)$ at $n$ do
6:   Compute for each $o\in\mathcal{O}$:
$\mathrm{UCB}(n,o,p)=\frac{V(n,o,p)}{N(n,o,p)+\epsilon}+c\;\sqrt{\frac{\ln\bigl{(}\sum_{o^{\prime}}N(n,o^{\prime},p)+1\bigr{)}}{\,N(n,o,p)+\epsilon\,}}$
7:   $o^{*}\leftarrow\arg\max_{o}\mathrm{UCB}(n,o,p)$
8:   if child $n^{\prime}$ annotated by $(o^{*},p)$ exists then
9:    $n\leftarrow n^{\prime},\quad p\leftarrow\neg p$
10:   else
11:    break
12:   end if
13:  end while2. EXPANSION:
14:  if no child of $n$ annotated by $(o^{*},p)$ then
15:   $\pi_{p}(n^{\prime})\leftarrow\textsc{LLMGenerate}\bigl{(}\pi_{p}(n),o^{*},p\bigr{)}$
16:   $\pi_{\neg p}(n^{\prime})\leftarrow\pi_{\neg p}(n)\quad$ {copy opponent’s code}
17:   Annotate $n^{\prime}$ with creator $(o^{*},p)$
18:   Add $n^{\prime}$ as a child of $n$
19:  else
20:   $n^{\prime}\leftarrow$ existing child
21:  end if3. SIMULATION:
22:  $C_{p}\leftarrow\textsc{Evaluate}\bigl{(}\pi_{p}(n^{\prime})\bigr{)}$
23:  $I_{p}\leftarrow(C_{0}-C_{p})/|C_{0}|\,(\%),\quad I_{\neg p}\leftarrow(C_{0}-C_{\neg p}(n^{\prime}))/|C_{0}|\,(\%)$
24:  $Q_{p}\leftarrow\lambda\,\sigma(I_{p})\;+\;(1-\lambda)\,\sigma(I_{p}-I_{\neg p})$ 4. BACKPROPAGATION:
25:  for each ancestor $a$ from $n^{\prime}$ back to $n_{0}$ do
26:   Let $(o_{a},p_{a})$ be the annotation on the edge from $a$
27:   $N(a,o_{a},p_{a})\leftarrow N(a,o_{a},p_{a})+1$
28:   $V(a,o_{a},p_{a})\leftarrow V(a,o_{a},p_{a})+Q_{p}$
29:  end for
30:  Best Update:
31:  if $C_{p}<C^{*}_{p}$ then
32:   $C^{*}_{p}\leftarrow C_{p},\quad\pi^{*}_{p}\leftarrow\pi_{p}(n^{\prime})$
33:  end if
34:  $p\leftarrow\neg p\quad$ {switch to other player}
35:end for
36:Return: $\pi^{*}_{1},\;\pi^{*}_{2}$
@@ C.4 Second Round: System-aware Refinement
1:Input: Initial combination $\boldsymbol{\Pi}^{(0)}=(\pi^{(0)}_{1},\dots,\pi^{(0)}_{K})$, baseline cost $C_{0}$, number of iterations per strategy $T$
2: Initialize global best combination $\boldsymbol{\Pi}^{*}\leftarrow\boldsymbol{\Pi}^{(0)}$, cost $C^{*}\leftarrow C_{0}$
@@
4:  Fix baseline $\boldsymbol{\Pi}_{\text{base}}\leftarrow\boldsymbol{\Pi}^{*}$, $C_{\text{base}}\leftarrow C^{*}$
5:  Initialize player best costs: $C_{1}\leftarrow\infty,\quad C_{2}\leftarrow\infty$
6:  Initialize current player $p\leftarrow 1$, failure count $f\leftarrow 0$
7:  for $t=1$ to $T$ do
8:   if $f\geq 3$ then
9:    Reset $\pi_{k}^{(t)}\leftarrow\pi_{k}^{\text{base}}\quad$ {fallback to baseline}
10:    $f\leftarrow 0$
11:   else
12:    $\pi_{k}^{(t)}\leftarrow$ best-performing implementation of player $p$ for strategy $k$
13:   end if
14:   Form context combination: $\boldsymbol{\Pi}^{(t)}\leftarrow\boldsymbol{\Pi}_{\text{base}}\text{ with }\pi_{k}\mapsto\pi_{k}^{(t)}$
15:   Apply Operator: $\hat{\pi}_{k}^{(t)}\leftarrow\textsc{LLMGenerate}(\boldsymbol{\Pi}^{(t)},\,k,\,p)$
16:   $\boldsymbol{\Pi}^{\prime(t)}\leftarrow\boldsymbol{\Pi}^{(t)}\text{ with }\pi_{k}\mapsto\hat{\pi}_{k}^{(t)}$
17:   Evaluate: $C^{(t)}\leftarrow\textsc{Evaluate}(\boldsymbol{\Pi}^{\prime(t)})$
18:   Compute improvement: $I^{(t)}\leftarrow\dfrac{C_{\text{base}}-C^{(t)}}{|C_{\text{base}}|}\cdot 100$
19:   if $C^{(t)}<C_{p}$ then
20:    $C_{p}\leftarrow C^{(t)}$
21:   end if
22:   if $C^{(t)}<\infty$ and $I^{(t)}>-50$ then
23:    $f\leftarrow 0$
24:   else
25:    $f\leftarrow f+1$
26:   end if
27:   $p\leftarrow\neg p\quad$ {switch player}
28:  end for
29:  Let $C_{k}^{\text{win}}\leftarrow\min(C_{1},C_{2})$
30:  if $C_{k}^{\text{win}}<C^{*}$ then
31:   $C^{*}\leftarrow C_{k}^{\text{win}}$, $\boldsymbol{\Pi}^{*}\leftarrow$ winning combination for strategy $k$
32:  end if
33:end for
34:Return: $\boldsymbol{\Pi}^{*},\,C^{*}$

### v2

@@ A.4 Self-play and Adversarial Learning in LLMs
Recent work has shown that self-play—where language models act as interacting agents—can lead to significant gains in reasoning quality, robustness, and decision-making. Inspired by game-theoretic dynamics, these methods structure LLM interaction through alternating roles such as attacker–defender, proposer–critic, or buyer–seller, enabling emergent behavior that improves model output beyond static prompting. For instance, SPAG (cheng2025selfplayingadversariallanguagegame) frames reasoning as a two-player adversarial game in which an attacker introduces misleading arguments and a defender must debunk them, improving the model’s capacity for deception detection. Similarly, the CDG framework (wangimproving) trains a pair of prover and critic agents to iteratively expose and correct reasoning flaws, leading to more logically sound outputs.
Other work explores negotiation and cooperation in self-play. Fu et al. (fu2023improving) design an in-context feedback loop in which LLMs simulate negotiation roles (e.g., buyer vs. seller) and revise their proposals based on AI-generated critiques, resulting in improved utility and coherence in generated dialogues. These methods typically rely on prompt structuring, role conditioning, and iterative critique to create a feedback-rich interaction landscape—one that mirrors social or argumentative reasoning among humans.
@@
Adversarial learning in LLMs also intersects with verbal reinforcement learning and reflective prompting. For example, Reflexion (shinn2023reflexion) guides the model to reflect on its own past errors and learn through verbalized feedback, strengthening multi-step reasoning capabilities. Although these studies focus primarily on tasks like logic puzzles, negotiation, and QA, their results suggest that interactive multi-agent dynamics offer a powerful foundation for tasks that benefit from critique, revision, or competition. Taken together, these approaches illustrate how adversarial pressure and structured self-interaction can improve LLM outputs by introducing feedback loops, multiple perspectives, and dynamic behavior adjustment—offering a foundation for more sophisticated systems that go beyond static prompting or single-agent search.
@@ B.1 Benchmark Problems
The Capacitated Vehicle Routing Problem (CVRP) extends the TSP by introducing multiple vehicles that serve customers under capacity constraints. Let $G=(V,E)$ be a complete undirected graph where $V=\{v_{0},v_{1},\dots,v_{n}\}$ consists of a depot node $v_{0}$ and $n$ customer nodes, and let $c(i,j)\in\mathbb{R}_{\geq 0}$ denote the travel cost between nodes $v_{i}$ and $v_{j}$. Each customer $v_{i}$ has a demand $d_{i}>0$, and all vehicles are identical with a fixed capacity $Q$. A route is a sequence $r=(v_{0},v_{i_{1}},\dots,v_{i_{m}},v_{0})$ in which a single vehicle departs from the depot, serves a subset of customers, and returns to the depot. The cost of such a route is given by:
@@ B.2 Benchmark Algorithms
Guided Local Search (GLS) is a well-established metaheuristic designed to escape local optima by penalizing frequently occurring features in poor-quality solutions. Introduced by VOUDOURIS1999469, GLS enhances standard local search by augmenting the objective function with feature-based penalties, thereby encouraging diversification while maintaining efficiency. It has demonstrated strong performance on various combinatorial problems, particularly the TSP, and has been used in hybrid solvers such as GLS-PR for Vehicle Routing Problems (VRP) (shaw1997new; shaw1998using) and more recently in neural-guided contexts (hudson2022graphneuralnetworkguided; sui2024neuralgls).
@@
2:Main strategy: $G\leftarrow\textsc{BuildGuideMatrix}(D)$
@@
4:for $t=1..T$ do
5:  for $m=1..M$ do
6:   $e\leftarrow\textsc{SelectEdgeToPenalize}(\sigma,G,P)$
7:   $P\leftarrow\textsc{Penalize}(P,e)$
8:   $\sigma\leftarrow\textsc{ImproveLocal}(\sigma,\textsc{AugmentCosts}(D,P))$
9:  end for
10:  $\sigma\leftarrow\textsc{ImproveLocal}(\sigma,D)$
@@
12:end for
13:Return $\sigma_{\text{best}}$
@@
Ant Colony Optimization (ACO) is a metaheuristic inspired by the foraging behavior of real ants, first introduced by dorigo1996ant as the Ant System for solving COP s. In ACO, a colony of artificial ants incrementally construct solutions by moving on a problem graph, guided by probabilistic rules that balance pheromone intensity and heuristic information. After completing a solution, each ant deposits pheromone on visited components, reinforcing promising paths over iterations. The Ant Colony System (ACS) (dorigo1997ant) refined this process by introducing local pheromone updates and elitist reinforcement strategies.
Since its inception, ACO has been extended in various directions, including hybridization with local search, parallel implementations, and learning-based variants. Recent approaches such as DeepACO (ye2023deepaco) incorporate neural networks to guide construction policies or predict pheromone distributions, significantly improving scalability and adaptability to different problem instances. These developments have established ACO as a versatile and extensible baseline for many routing and packing tasks.
@@
We now formalize the ACO solver structure used across our benchmark tasks. Each solver consists of three strategies—initialization, transition rule, and pheromone update—optimized independently. We provide a unified pseudocode and detail problem-specific implementations in Table 5. While prior work in AHD typically focuses on optimizing a single heuristic matrix $H$ (e.g., based on inverse distances), our formulation broadens this space in several ways. First, we additionally treat the initialization of the pheromone matrix $P$ as part of the optimization process. Second, we model the combination of $H$ and $P$ into transition weights as a learnable strategy, rather than using a fixed formula. Third, we expose the pheromone update rule itself to adaptation. Finally, we extend the function signatures to include both the current iteration $t$ and the total horizon $T$, enabling LLMs to discover temporally adaptive behaviors that prior formulations cannot express.
@@
The idea of deconstruction and repair is not new in combinatorial optimization. Hottung_Andr__2020 proposed a neural variant of Large Neighborhood Search (LNS), where destroy and repair policies are jointly learned via reinforcement learning. wu2021learninglargeneighborhoodsearch further improved this framework by leveraging GNN-based representations for adaptive neighborhood selection. More recently, fu2023hierarchicaldestroyrepairapproach introduced a hierarchical destroy-and-repair framework that scales to million-city TSP instances by progressively refining solutions across multiple levels. In parallel, li2025destroyrepairusinghyper explored hypergraph-based operators to guide destruction and reconstruction phases, achieving state-of-the-art results.
@@ C.1 Prompt Structure
The system prompt is a persistent instruction that defines the identity and behavior of the lLLM across interactions. Unlike user inputs, which vary per query, it provides a stable context that aligns the model’s responses with task objectives. A common technique is role-based prompting, where assigning the LLM a specific persona—e.g., “competitive algorithm designer”—helps activate relevant reasoning styles. This role-play approach has been shown to enhance zero-shot performance by encouraging coherence, structured thinking, and task relevance (kong2024betterzeroshotreasoningroleplay).
@@ 3. Ensure code is syntactically correct and handles edge cases
Implement the initialization strategy that sets up guidance matrices for route finding:. 1import numpy as np 2 3def initialize(distances: np.ndarray) -> tuple[np.ndarray, np.ndarray]: 4 """ 5 Initialize heuristic and pheromone matrices for route optimization. 6 7 distances : np.ndarray, shape (n_cities, n_cities) 8 heuristic : np.ndarray, shape (n_cities, n_cities) 9 Matrix representing desirability of traveling between cities 10 pheromone : np.ndarray, shape (n_cities, n_cities) 11 Matrix representing initial intensity of guidance trails 12 """ 13 # Your implementation here 14 pass <</task>>
@@
The human message, framed from a third-person referee perspective, delivers the competition status to the agent while simultaneously acting as a coach—offering guidance and exerting competitive pressure. Notably, we intentionally place the operator instructions at the end of the prompt to draw the model’s attention, leveraging findings from liu2023lostmiddlelanguagemodels, which show that LLMs often overlook information positioned in the middle of long prompts.
@@ C.3 First Round: Component-wise Competition
1:Input: initial strategies $\boldsymbol{\Pi}$, outer iters $T_{\text{outer}}$, inner iters $T_{\text{inner}}$
2: Evaluate baseline $C_{0}\leftarrow\textsc{Evaluate}(\boldsymbol{\Pi})$
3: Initialize trees $\{\mathcal{T}_{k}\}$ and stats $\{N_{k},R_{k}\}$ for each strategy
4:for $t=1$ to $T_{\text{outer}}$ do
5:  Select promising strategy $k^{\ast}$ via UCB rule
@@
7:  Evaluate updated strategy set $\boldsymbol{\Pi}$
8:  if improvement over baseline then
9:   Update reward $R_{k^{\ast}}$ and baseline $C_{0}$
10:   Propagate new baseline to all trees
11:  else
12:   Apply small reward and revert strategy
13:  end if
14:  Increment visit count $N_{k^{\ast}}$
15:end for
16:Return: Final strategies $\boldsymbol{\Pi}$
@@
where $F(s\!\to\!s^{\prime})=U(s^{\prime})-U(s)$ is exactly in the form of a potential‐based shaping reward ($\Phi(s^{\prime})-\Phi(s)$), which by ng1999policy guarantees policy invariance. $G(s)=U(s)$ depends only on the current state $s$ (i.e. is constant across all child actions), and thus does not affect the relative ranking of successor Q‐values in MCTS selection.
@@
1:Input: root node $n_{0}$ with strategies $(\pi_{1},\pi_{2})$, baseline $C_{0}$, operator set $\mathcal{O}$, iterations $T$
2: Initialize best costs $C^{*}_{1},C^{*}_{2}$ and best strategies $\pi^{*}_{1},\pi^{*}_{2}$
@@
4:  $n\leftarrow n_{0}$, set current player $p\leftarrow 1$
5:  (1) Selection: Traverse tree by choosing operator $o^{*}$ with highest UCB until an expandable node is found.
6:  (2) Expansion: Generate a new child $n^{\prime}$ by applying operator $o^{*}$ on $\pi_{p}(n)$, while copying opponent’s code.
7:  (3) Simulation: Evaluate new strategies, compute improvement scores, and obtain quality $Q_{p}$.
8:  (4) Backpropagation: Update visit counts and accumulated values along the path to the root. If $C_{p}$ improves, record as new best $\pi^{*}_{p}$ and $C^{*}_{p}$.
9:  Switch player $p\leftarrow\neg p$
10:end for
11:Return: Best strategies $\pi^{*}_{1},\;\pi^{*}_{2}$
@@ C.4 Second Round: System-aware Refinement
1:Input: initial strategy set $\boldsymbol{\Pi}^{(0)}$, baseline cost $C_{0}$, iterations per strategy $T$
2: Initialize global best $(\boldsymbol{\Pi}^{*},C^{*})\leftarrow(\boldsymbol{\Pi}^{(0)},C_{0})$
@@
4:  Set baseline $(\boldsymbol{\Pi}_{\text{base}},C_{\text{base}})\leftarrow(\boldsymbol{\Pi}^{*},C^{*})$
5:  Initialize per-player best costs $C_{1},C_{2}$ and current player $p\leftarrow 1$
6:  for $t=1$ to $T$ do
7:   Select current implementation of player $p$ for strategy $k$
8:   Form context combination by replacing $\pi_{k}$ in $\boldsymbol{\Pi}_{\text{base}}$
9:   Generate candidate $\hat{\pi}_{k}$ using LLM
10:   Evaluate new combination to obtain cost $C^{(t)}$
11:   Update player’s best if $C^{(t)}$ improves
12:   Handle failure or fallback if no improvement
13:   Switch player $p\leftarrow\neg p$
14:  end for
15:  Determine winner between players and update global best $(\boldsymbol{\Pi}^{*},C^{*})$ if improved
16:end for
17:Return: Final optimized set $\boldsymbol{\Pi}^{*},\,C^{*}$